# DiLi: A Lock-Free Asynchronously Distributable Linked List

### v1

@@
Abstract
Modern databases use dynamic search structures that store a huge amount of data, and often serve them using multi-threaded algorithms to support the ever-increasing throughput needs. When this throughput need exceeds the capacity of the machine hosting the structure, one either needs to replace the underlying hardware (an option that is typically not viable and introduces a long down time) or make the data structure distributed. Static partitioning of the data structure for distribution is not desirable, as it is prone to uneven load distribution over time, and having to change the partitioning scheme later will require downtime.
Since a distributed data structure, inherently, relies on communication support from the network stack and operating systems, we introduce the notion of conditional lock-freedom that extends the notion of lock-free computation with reasonable assumptions about communication between processes. We present DiLi, a conditional lock-free, linearizable, and distributable linked list that can be asynchronously and dynamically (1) partitioned into multiple sublists and (2) load balanced by distributing sublists across multiple machines. DiLi contains primitives for these that also maintain the lock-free property of the underlying search structure that supports find, remove, and insert of a key as the client operations.
Searching for an item in DiLi is by a novel traversal that involves a binary search on the partitioning scheme, and then a linear traversal on a limitable number of linked nodes. As a result, we are able to empirically show that DiLi performs as well as the state-of-the-art lock-free concurrent search structures that are based off of a linked list when executed on a single-machine. We also show that the throughput of DiLi scales linearly with the number of machines that host it.
keywords:
@@ Abstract
[…]
Eligible for best student paper award.
1 Introduction
Concurrent dynamic search structures play a crucial role in many modern databases [3, 24, 12]. With the rising number of concurrent operations to a database, memory contention should be managed with care to preserve the atomicity of its operations. For this, locks are often utilized, which can lead to pitfalls such as deadlock, priority inversion, and convoying [18].
Lock-free search structures avoid the drawbacks of lock-based methods but are limited to single-machine shared memory systems. Scaling further will require either a hardware upgrade to the single machine and perform a data migration, or a scheme to distribute load across multiple machines, with each machine handling only a ‘part’ of the data structure. The former can be expensive if the machine is self-hosted (for privacy reasons), or can be limited to the highest available instance if cloud-hosted. Migrations here would still cause downtime to the database. Further, the need to distribute also arises from the observation that Moore’s law now has diminishing returns [22].
Partitioning for load balancing can be static (fixed key distribution per machine) or dynamic (key distribution evolves over time). Static schemes cannot adapt to workload shifts or scale without downtime. Dynamic schemes offer flexibility, but require shared memory structures to be temporarily unavailable during redistribution. It is, therefore, necessary to develop dynamic partitioning that supports redistribution asynchronously, i.e., while preserving the properties of the client operations.
In this paper, we focus on conditional lock-free linked lists. Linked lists are fundamental data structures that can in turn construct other data structures [30]. Specifically, we distribute the well-known Harris Linked list [14], and improve its performance even on a single-machine setup. Due to the distributed scalability and the general performance improvement, any multi-threaded lock-free data structure that is implemented with Harris List as its foundation, can automatically become a faster and distributable data structure by replacing Harris List with DiLi for its foundation.
Contributions: We introduce a new distributed linked list DiLi – that performs find, insert and remove of a key both in a multi-threaded and a multi-machine setup while supporting asynchronous dynamic partitioning. Any part of this data structure (henceforth called a sublist) can be split into two smaller parts. The resulting partitioning scheme additionally serves as an index for quick search traversals, resulting in a significant performance boost that is sufficient to become comparable to lock-free skip lists [11] in workloads of varying read proportions (as we show empirically).
@@ 1 Introduction
Empirically, we show that 1. The performance of DiLi is comparable to that of a lock-free skip list in multiple workloads, 2. DiLi is capable of automatic scaling upto several machines with a linear growth in throughput, and 3. The performance of Split and Move operations are practical even under write-intensive workloads.
Outline: We introduce the notion of conditional lock-freedom in Section˜2, introduce our system model in Section˜3, and explain our design principles in Section˜4. DiLi is then described through the following three sections: Algorithm (Section˜5), Correctness (Section˜6) and Empirical Evaluation (Section˜7). Then we compare DiLi with related work in Section˜8. Finally, the concluding remarks are in Section˜9.
2 Extending Lock-Freedom for Distributed Systems
Even in simple client-server models such as a single server-client model, the traditional lock-freedom definition: “A method is lock-free if it guarantees that infinitely often some method call finishes in a finite number of steps” [18] is difficult to achieve, as any client request has to get past unbounded network delays to get the operation request on a data structure served. Also, the client thread that initiates this request typically goes to sleep (hence having a dependency on another thread) until it receives a response from an associated server machine thread. In other words, the definition of traditional lock freedom cannot be satisfied in any client-server machine (or more generally in distributed systems). Due to the above constraints, we condition any communication between threads (such as messages) in the distributed system to only take a finite number of steps (e.g., with reliable channels).
@@ 2 Extending Lock-Freedom for Distributed Systems
3 System Model
We study a list sorted by key and refer to each of its elements as elements. To construct our partitioned version of this list, we introduce the notion of a sublist that comprises a set of items contained by a subhead and a subtail. For example, in Figure˜1, we show a list consisting of 3 sublists (SL1 to SL3), each with a subhead (SH) and a subtail (ST) that hold a range of keys. The list supports the following three operations: 1. Find(key): Returns false if the key is not found in the list, and otherwise returns true. 2. Insert(key): Returns false if the key is already in the list and otherwise inserts the key in its sorted position, returning true. 3. Remove(key): Returns false if the key is not found in the list, and otherwise deletes the key, de-links the item off the list, and returns true. Additionally, a server may invoke operations Split (to partition the submit into two sublists), Move (to create a live clone of the sublist to another server), and Switch (to transfer ownership of sublist to new server). [^1] We also support a Merge operation for two adjacent sublists. Due to the similarity of its logic to the Split algorithm, we have pushed it to Appendix B.
Figure 1: A sample list for keys 1–100 being composed of a set of 3 sublists.
@@ 3 System Model
As shown in Figure˜2, we follow a decentralized architecture. Each client will be assigned a specific server X (X is any server in the system) to invoke its list operations for any key $k$ through RPCs. The server X responds to the client, by invoking RPC to server Y that is capable of performing the list operation for $k$. In this case, X acted as a proxy and is said to have delegateddelegated the request to Y. Note that X could also receive delegated requests from other servers as it also owns some sublists. Naturally, in a single machine setup, this overhead is avoided, as X will always be the same as Y. Further, we assume that a server machine does not crash entirely, and has atleast one usable core. This is essential to serve the keyset that the machine is responsible for, as we do not create replicas of the list.
An important concern when aiming for linearizability in a distributed system is to operate under the constraints of the CAP theorem [7]. By ensuring that any operation takes place on only one sublist, and that a sublist entirely resides in one single machine at any given time, we are able to overcome the effects of network partitioning that impedes linearizability.
@@
Figure 2: Client invoking an RPC request on a key to its assigned Server X and server X delegating the request to server Y that owns the sublist for that key.
4 Design Principles
The design of the distributed list involves several techniques in order to provide a linearizable lock-free implementation. The first is a smart pointer, which is inspired by the Harris Linked List [14]. Harris used the least significant bit (LSB) of a list item’s next pointer to denote whether an item has been soft deleted, as it is otherwise always set to 0 due to address alignment. In doing so, the time point of checking the value of the next pointer became crucial for linearizability arguments. In addition, there is common limitation on the x86-64 MMUs of commodity hardware (such as without 5-level paging extensions [1]) only to have a 48-bit virtual address space. Due to this, the 16 most significant bits of the 64-bit pointer remain unused during memory allocations. DiLi uses these 16 bits to globally indicate the server ID that the distributed list item belongs to.
Secondly, we maintain all sublists as entries in a lazily replicated registry—an index where each entry has a smart subhead pointer that encodes the server ID. This ID identifies the server that hosts the sublist and guides request routing. Although lazy replication may cause temporary staleness and latency, safeguards (see Section˜5) ensure that lock-freedom and linearizability remain intact. Lazy replication allows the registry to remain sorted via copy-on-write inserts. This enables logarithmic-time key lookup. This leads to our next design principle – the hybrid search where we use a logarithmic search in the registry followed by a linear search in the sublist. By keeping a check on the number of items present in a sublist, we achieve very quick traversals in the distributed list.
@@ 4 Design Principles
The move terminates, activates the cloned sublist, and transfers ownership to the new server when both machines have identical copies of the sublist. Although blocking client operations simplifies this, our approach achieves it asynchronously using integer counters around sublist updates. The Switch operation then needs to update the registry in all servers before another Move can happen on this sublist. In the interim, careful rerouting (called delegations) of requests should take place as the registry can be temporarily out-dated.
5 DiLi: A Distributable List
@@
DiLi is a linearizable sorted linked list with a two-layered data structure (described in Section˜5.1) – (1) an index called the Registry, that book keeps the subhead and subtail pointers (to shortcut into a sublist), and (2) sublists that consist of the underlying list items. Upon reaching a subhead, a linear traversal is employed to find the search node. The find, insert, and remove operations (as described in Section˜5.2) that utilize this traversal are lock-free, but only operate on the sublist space. For ease of explanation, we use colors in the algorithms. Black lines are executed the same way, regardless of whether there is an ongoing background operation, the red lines adapt to a concurrent Split and hence will be explained as part of Section˜5.3, and the blue lines adapt to a concurrent Move/Switch and are explained in Section˜5.4. The novel contribution in our approach is an asynchronous yet versatile Split, Move, and Switch algorithms that can be used to build any partitioning and load balancing scheme of choice to suit the workload of interest.
5.1 Data Structures: List Item and Registry
@@
Algorithm˜1 depict the two building blocks of DiLi. Like any list item, DiLi items contain a key to search for the item and a next pointer to lead to the next item. Note that this is a distributed list, with a traversal guided by the next pointer. We denote the type of this pointer as Ref. Any Ref pointer, say X, needs to be operated on by a bit mask to force the first 16 bits and the last bit to 0 before accessing the address that X points to. We overload this functionality into the C++ pointer access notation ($\rightarrow$). Fetching serverID bits, and if the item has been marked, are also implemented through bitmasks. For convenience, we denote them as X.id and X.mark in the algorithms.
@@
The registry consists of sublist entries and supports getByKey(key) operation to find the sublist entry to which a searched key belongs. Operation addEntry(entry) is used to insert an entry into an already sorted registry. We utilize a copy-on-write (COW) technique to update the registry pointer, to atomically update the registry with a new sorted registry with an added entry. Because registry replication is done lazily, it is sufficient to have only one update happening at a time on the registry. Hence any balanced single-writer tree data structure (binary tree, B-tree, or even a skip list) could also be used to build the registry, thereby offering a logarithmic search for getByKey(key) operation. For the sake of simplicity, we maintain the registry as a binary search tree in a large array and keep track of its size. We also utilize a variant of hazard pointers [25] to reclaim memory when an old pointer of the registry (after overwritten by COW) is no longer referenced by any thread. To save space, the functioning of the registry is explained in Appendix˜A.
1
2struct Item:
3  Key key;
4  Key keyMax;
5  int ts;
6  int sId;
7  Ref next;
8  int stCt ;
9  int endCt ;
10  Ref newLoc ;
11
12struct Entry:
13  Ref subhead;
14  Ref subtail;
15  Key keyMin;
16  Key keyMax;
17  int stCt ;
18  int endCt ;
19  int offset;
20
21struct Registry:
22  Entry * entries[MAX_SUBLISTS] = $\{null\cdots null\}$;
23  int size = 0;
24
25
Algorithm 1 Data Structures: Item and Registry
5.2 Client Operations
@@
All client operations perform the following in order: 1. A binary search on the replicated registry to retrieve the sublist entry for a key. 2. Routing the request to the appropriate server based on the server ID of the subhead present in the sublist entry. 3. A linear traversal on the sublist to perform operations in a Harris-List fashion with retries beginning from the subhead of the sublist.
@@
The end machine that serves the sublist traverses using the Search algorithm in Algorithm˜2. It resembles Harris List search, where the left and right nodes to the searched key are returned, while delinking any marked items encountered along the traversal. However, we optimize this search by limiting the search to start from the pre-determined subhead, instead of the list head. In our case, the rightNode is the item of the searched key or the one immediately greater present in the list. leftNode is the corresponding previous item.
Find and Remove utilize the returned right node of a search to determine if a key is present. Find returns its response based on this check (after Line 2), while Remove continues with a Delete on the found item using the Delete function. The Delete attempts to soft delete the node by marking its next pointer’s mark bit using CAS. If the CAS succeeds (Line 2), then the Remove returns true, and if it is already marked, then it returns false.
The Insert operation as described in Algorithm˜3 utilizes the InsertInSublist routine after determining the sublist to insert the key. This routine begins with a search operation to determine the left and right node window. If the right node matches the key to be inserted, then the Insert returns false. If the next pointer of left node is altered by getting marked or through a competing insert, then the CAS fails and the search is restarted on the same sublist to retry the insert. The logic behind the stCt, endCt, newLoc and ts assignments become apparent only during the background operations, as covered in the next sections.
26
27Ref Search(Key key, Ref head, Ref &leftNode):
28  curr = head;
29
@@
32
33  while true do
34     leftNode = prev = curr;
35     $curr\rightarrownext$;
@@
38        if head.id == me
39           leftNode = prev = head;
@@
43
44          continue;
45
46       return head;
47
@@
50           return null;
51
52        $curr\rightarrownext$;
53        if curr.id $\neq$ me
54           return curr;
55
56       leftNode = prev = curr;
57        head = curr;
58        $curr\rightarrownext$;
59
@@ 58        $curr\rightarrownext$;
61        if delinkNode(prev, curr)
62           Lines 2 to 2;
63           continue;
64
65       else
66           leftNode == prev == head;
@@ 62           Lines 2 to 2;
81
82  return null;
83
84bool delinkNode(Ref prev, Ref &curr):
@@
86  while currNext.mark = 1 do
@@
88
89 temp = curr;
90  if CAS(prev,temp,currNext)
91     curr = currNext;
92     return true;
93
94 return false;
95
96bool Find(Key key, Ref SH = null):
@@
98     entry = registry.getByKey(key);
@@
100
101 sId = SH.id;
102  if sId == me
103     leftNode = null;
104     node = Search(key, SH, leftNode);
105     if node == null
106        return false;
107
108    sId = node.id;
109     if sId == me
110        return true;
111
112     response = Send Find(key,node) to sId;
113     return response;
114
115 response = Send Find(key,SH) to sId;
116  return response;
117
118bool Remove(Key key, Ref SH = null):
119  Same as Find but line 2 is changed to
120 return Delete(node, key);
121  and the corresponding delegates are that of Remove instead of Find.
122bool Delete(Ref node, Key key):
123  result = false;
@@
125     return false;
126
[…]
128
[…]
131     return response;
132
133  while true do
[…]
136        break;
137
[…]
139     newVal = temp;
140     newVal.mark = true;
[…]
142        result = true;
[…]
145
146       else
[…]
148
149       break;
150
151 leftNode = null;
[…]
153  Search(key, SH, leftNode);
154  return result;
155
Algorithm 2 Distributed List Traversal, Find and Remove
156
@@
158     entry = registry.getByKey(key);
@@
160
161 sId = SH.id;
162  if sId == me
163     return InsertInSublist(key, SH);
164
165 response = Send Insert(key,SH) to sId;
166  return response;
167
168Entry *Split(Entry * entry, Ref sItem):
169  *newstCt = new int[1];
170  *newendCt = new int[1];
171  SH = new Item(SH_KEY, $\phi$, ts.fetch_add(1), me, null, newstCt, newendCt, null);
[…]
173
174 do
[…]
176     if temp.mark
177        return null;
178
[…]
181
[…]
184  do
185     prev = curr;
[…]
188     $curr\rightarrownext$;
189
[…]
191 do
192
193    $a_{1}$ = newstCt- newendCt ;
194     $a_{2}$ = $\rightarrowstCt- sItem\rightarrowendCt$ ;
195
196 while $a_{1}+a_{2}\neq entry\rightarrow offset$ ;
[…]
198
199 registry.addEntry(newEntry);
@@
206
207 return newEntry;
208
209bool RegisterSublistRecv(Key keyMin, Ref SH):
210  Entry leftEntry = registry.getByKey(keyMin);
@@
[…]
213  return true;
214
215bool InsertInSublist(Key key, Ref SH):
216  while true do
217     leftNode = null;
218     rightNode = Search(key, SH, leftNode);
219     if rightNode $\neq$ null
220        if rightNode.id == me
221           return false;
222
223       response = Send Insert(key,rightNode) to rightNode.id;
224        return response;
225
@@
229           return false;
230
[…]
232        if SH.id == me
233           continue;
234
[…]
236        return response;
237

### v2

# Transforming Lock-free Linked Lists into Distributed Lock-free Linked Lists

@@
Abstract.
Modern databases use dynamic search structures that store an enormous amount of data, and often serve them using multi-threaded algorithms to support the ever-increasing throughput needs. When this throughput need exceeds the capacity of the machine hosting the structure, one either needs to replace the underlying hardware (an option that is typically not viable and introduces a long down time) or make the data structure distributed. Static partitioning of the data structure for distribution is not desirable, as it is prone to uneven load distribution over time, and having to change the partitioning scheme later will require downtime.
The goal of this paper is to extend a concurrent data structure to distributed data structures that provide dynamic load balancing while preserving important properties such as lock freedom. With this intuition, first, we introduce the notion of conditional lock-freedom that extends the notion of lock-free computation with reasonable assumptions about communication between processes. Then, we present DiLi, a conditional lock-free, linearizable, and distributable linked list that can be asynchronously and dynamically (1) partitioned into multiple sublists and (2) load balanced by distributing sublists across multiple machines. DiLi contains primitives for these that also maintain the lock-free property of the underlying search structure that supports find, remove, and insert of a key as the client operations.
We show that DiLi bridges the gap between concurrent data structures and distributed data structures. Specifically, DiLi provides comparable (and better in write-intensive workloads) performance to skip lists (which are typically the fastest data structures for search in a concurrent environment). Plus, it provides horizontal scaling by permitting dynamic load balancing on a distributed environment.
@@ Abstract.
$c$s: Theory of computation Data structures design and analysisccs: Computing methodologies Distributed computing methodologies Distributed algorithms
1. Introduction
Concurrent dynamic search structures play a crucial role in many modern databases (Apache, 2008, 2012, 2011; Prout et al., 2022). With the rising number of concurrent operations to a database, memory contention should be managed with care to preserve the atomicity of its operations. For this, locks are often utilized, which can lead to pitfalls such as deadlock, priority inversion, and convoying (Herlihy and Shavit, 2012).
Lock-free search structures avoid the drawbacks of lock-based methods but are limited to single-machine shared memory systems. Scaling further will require either a hardware upgrade to the single machine and perform a data migration, or a scheme to distribute load across multiple machines, with each machine handling only a ‘part’ of the data structure. The former can be expensive if the machine is self-hosted (for privacy reasons), or can be limited to the highest available instance if cloud-hosted. Migrations here would still cause downtime to the database. Further, the need to distribute also arises from the observation that Moore’s law now has diminishing returns (Leiserson et al., 2020).
Partitioning for load balancing can be static (fixed key distribution per machine) or dynamic (key distribution evolves over time). Static schemes cannot adapt to workload shifts or scale without downtime. Dynamic schemes offer flexibility, but require shared memory structures to be temporarily unavailable during redistribution. With this intuition, we consider the question: Is it possible to partition and dynamically load balance a data structure while preserving the property of lock freedom. As an instance, we focus on the list data structure, Specifically, we focus on conditional lock-free (that extends lock-freedom to distributed environments) linked lists. Linked lists are fundamental data structures that can in turn construct other data structures (Valois, 1995). Specifically, we distribute the well-known Harris Linked list (Harris, 2001), and improve its performance even on a single-machine setup. Due to the distributed scalability and the general performance improvement, any multi-threaded lock-free data structure that is implemented with Harris List as its foundation, can automatically become a faster and distributable data structure by replacing Harris List with DiLi for its foundation.
Contributions: We introduce a new distributed linked list DiLi – that performs find, insert and remove of a key both in a multi-threaded and a multi-machine setup while supporting asynchronous dynamic partitioning. Any part of this data structure (henceforth called a sublist) can be split into two smaller parts.
@@ 1. Introduction
We design DiLi for write-heavy workloads and as heavily investigated in synchrobench (Gramoli, 2015), lock-free skip lists outperform lock-free trees in write-heavy workloads due to the lack of rebalancing overhead. Hence, we compare DiLi with a popular lock-free skip list(Fraser, 2004) (all references to skip list in this paper refer to (Fraser, 2004)). Empirically, we show that (1) The performance of DiLi is not only comparable to lock-free skip lists, but even faster in write heavy workloads. (2) DiLi is capable of automatic scaling upto several machines with a linear growth in throughput, and (3) The performance of Split and Move operations are practical even under write-intensive workloads.
Outline: We introduce the notion of conditional lock-freedom in Section 2, introduce our system model in Section 3, and explain our design principles in Section 4. DiLi is then described through the following three sections: Algorithm (Section 5), Correctness (Section 6) and Empirical Evaluation (Section 7). Then we compare DiLi with related work in Section 8. Finally, the concluding remarks are in Section 9.
2. Extending Lock-Freedom for Distributed Systems
Even in simple client-server models such as a single server-client model, the traditional lock-freedom definition: “A method is lock-free if it guarantees that infinitely often some method call finishes in a finite number of steps” (Herlihy and Shavit, 2012) is difficult to achieve, as any client request has to get past unbounded network delays to get the operation request on a data structure served. Also, the client thread that initiates this request typically sleeps (hence having a dependency on another thread) until it receives a response from an associated server machine thread. In other words, the definition of traditional lock freedom cannot be satisfied in any client-server machine (or more generally in distributed systems). Due to the above constraints, we condition any communication between threads (such as messages) in the distributed system to only take a finite number of steps (e.g., with reliable channels).
@@ 2. Extending Lock-Freedom for Distributed Systems
3. System Model
We study a list sorted by key and refer to each of its elements as items. To construct our partitioned version of this list, we introduce the notion of a sublist that comprises a set of items contained by a subhead and a subtail. For example, in Figure 1, we show a list consisting of 3 sublists (SL1 to SL3), each with a subhead (SH) and a subtail (ST) that hold a range of keys. The list supports the following three operations: (1) Find(key): Returns false if the key is not found in the list, and otherwise returns true. (2) Insert(key): Returns false if the key is already in the list and otherwise inserts the key in its sorted position, returning true. (3) Remove(key): Returns false if the key is not found in the list, and otherwise deletes the key, de-links the item off the list, and returns true. Additionally, a server may invoke operations Split (to partition the sublist into two sublists), Move (to create a live clone of the sublist in another server), and Switch (to transfer ownership of a sublist to a new server). We also support a Merge operation for two adjacent sublists. Due to the similarity of its logic to the Split algorithm, we have pushed it to Appendix(Me, 2025b).
Figure 1. A sample list for keys 1–100 being composed of a set of 3 sublists.
@@ 3. System Model
As shown in Figure 2, we follow a decentralized architecture. Each client will be assigned a specific server X (X is any server in the system) to invoke its list operations for any key $k$ through RPCs. The server X responds to the client, by invoking RPC to server Y that is capable of performing the list operation for $k$. In this case, X acted as a proxy and is said to have delegateddelegated the request to Y. Note that X could also receive delegated requests from other servers as it also owns some sublists. Naturally, in a single machine setup, this overhead is avoided, as X will always be the same as Y. Further, we assume that a server machine does not crash entirely, and has atleast one usable core. This is essential to serve the keyset that the machine is responsible for, as we do not create replicas of the list.
An important concern when aiming for linearizability(Herlihy and Wing, 1990) in a distributed system is to operate under the constraints of the CAP theorem (Brewer, 2000). By ensuring that any operation takes place on only one sublist, and that a sublist entirely resides in one single machine at any given time, we are able to overcome the effects of network partitioning that impedes linearizability.
@@
Figure 2. Client invokes RPC on a key to its assigned Server X and X delegates the request to Server Y that owns the key.
4. Design Principles
The design of the distributed list involves several techniques in order to provide a linearizable lock-free implementation. The first is a smart pointer, which is inspired by the Harris Linked List (Harris, 2001). Harris used the least significant bit (LSB) of a list item’s next pointer to denote whether an item has been soft deleted, as it is otherwise always set to 0 due to address alignment. In doing so, the time point of checking the value of the next pointer became crucial for linearizability arguments. In addition, there is common limitation on the x86-64 MMUs of commodity hardware (such as without 5-level paging extensions (1)) only to have a 48-bit virtual address space. Due to this, the 16 most significant bits of the 64-bit pointer remain unused during memory allocations. DiLi uses these 16 bits to globally indicate the server ID that the distributed list item belongs to.
Secondly, we maintain all sublists as entries in a lazily replicated registry—an index where each entry has a smart subhead pointer that encodes the server ID. This ID identifies the server that hosts the sublist and guides request routing. Although lazy replication may cause temporary staleness and latency, safeguards (see Section 5) ensure that lock-freedom and linearizability remain intact. Lazy replication allows the registry to remain sorted via copy-on-write inserts. This enables logarithmic-time key lookup. This leads to our next design principle – the hybrid search where we use a logarithmic search in the registry followed by a linear search in the sublist. By keeping a check on the number of items present in a sublist, we achieve very quick traversals in the distributed list.
@@ 4. Design Principles
The move terminates, activates the cloned sublist, and transfers ownership to the new server when both machines have identical copies of the sublist. Although blocking client operations can simplify this, our approach achieves it asynchronously using integer counters around sublist updates. The Switch operation then needs to update the registry in all servers before another Move can happen on this sublist. In the interim, careful rerouting (called delegations) of requests should take place as the registry can be temporarily out-dated.
5. DiLi: A Distributable List
@@
DiLi is a linearizable sorted linked list with a two-layered data structure (described in Section 5.1) – (1) an index called the Registry, that book keeps the subhead and subtail pointers (to shortcut into a sublist), and (2) sublists that consist of the underlying list items. Upon reaching a subhead, a linear traversal is employed to find the search node. The find, insert, and remove operations (as described in Section 5.2) that utilize this traversal are lock-free, but only operate on the sublist space. For ease of explanation, we use colors in the algorithms. Black lines are executed the same way, regardless of whether there is an ongoing background operation, the red lines adapt to a concurrent Split and hence will be explained as part of Section 5.3, and the blue lines adapt to a concurrent Move/Switch and are explained in Section 5.4. The novel contribution in our approach is an asynchronous yet versatile Split, Move, and Switch algorithms that can be used to build any partitioning and load balancing scheme of choice to suit the workload of interest.
5.1. Data Structures: List Item and Registry
@@
Algorithm 1 depict the two building blocks of DiLi. Like any list item, DiLi items contain a key to search for the item and a next pointer to lead to the next item. Note that this is a distributed list, with a traversal guided by the next pointer. We denote the type of this pointer as Ref. Any Ref pointer, say X, needs to be operated on by a bit mask to force the first 16 bits and the last bit to 0 before accessing the address that X points to. We overload this functionality into the C++ pointer access notation ($\rightarrow$). Fetching serverID bits, and if the item has been marked, are also implemented through bitmasks. For convenience, we denote them as X.id and X.mark in the algorithms.
@@
The registry consists of sublist entries and supports a getByKey(key) operation to find the sublist entry to which a searched key belongs. Operation addEntry(entry) is used to insert an entry into an already sorted registry. We utilize a copy-on-write (COW) technique to update the registry pointer, to atomically update the registry with a new sorted registry with an added entry. Because registry replication is done lazily, it is sufficient to have only one update happening at a time on the registry. Hence any balanced single-writer tree data structure (binary tree, B-tree, or even a skip list) could also be used to build the registry, thereby offering a logarithmic search for the getByKey(key) operation. For the sake of simplicity, we maintain the registry as a binary search tree in a large array and keep track of its size. We also utilize a variant of hazard pointers (Michael, 2004) to reclaim memory when an old pointer of the registry (after being overwritten by COW) is no longer referenced by any thread. To save space, the functioning of the registry is explained in Appendix(Me, 2025b). We also reserve two keys out of the available key range for the list to indicate the subhead (SH_KEY) and subtail (ST_KEY) items.
5.2. Client Operations
@@
All client operations perform the following in order: (1) A binary search on the replicated registry to retrieve the sublist entry for a key. (2) Routing the request to the appropriate server based on the server ID of the subhead present in the sublist entry. (3) A linear traversal on the sublist to perform operations in a Harris-List fashion with retries beginning from the subhead of the sublist.
@@
The end machine that serves the sublist traverses using the Search algorithm in Algorithm 1. It resembles Harris List search, where the left and right nodes to the searched key are returned, while delinking any marked items encountered along the traversal. However, we optimize this search by limiting the search to start from the pre-determined subhead, instead of the list head. In our case, the rightNode is the item of the searched key or the one immediately greater present in the list. leftNode is the corresponding previous item.
Find and Remove utilize the returned right node of a search to determine if a key is present. Find returns its response based on this check (after Line 1), while Remove continues with a Delete on the found item using the Delete function. The Delete attempts to soft delete the node by marking its next pointer’s mark bit using CAS. If the CAS succeeds (Line 1), then the Remove returns true, and if it is already marked, then it returns false.
The Insert operation as described in Algorithm 1 utilizes the InsertInSublist routine after determining the sublist to insert the key. This routine begins with a search operation to determine the left and right node window. If the right node matches the key to be inserted, then the Insert returns false. If the next pointer of left node is altered by getting marked or through a competing insert, then the CAS fails and the search is restarted on the same sublist to retry the insert. The logic behind the stCt, endCt, newLoc and ts assignments becomes apparent only during the background operations, as covered in the next sections.
struct Item:
  Key key;
  Key keyMax;
  int ts;
  int sId;
  Ref next;
  int stCt ;
  int endCt ;
  Ref newLoc ;
struct Entry:
  Ref subhead;
  Ref subtail;
  Key keyMin;
  Key keyMax;
  int stCt ;
  int endCt ;
  int offset;
struct Registry:
  Entry * entries[MAX_SUBLISTS] = $\{null\cdots null\}$;
  int size = 0;
Ref Search(Key key, Ref head, Ref &leftNode):
  curr = head;
@@
  end if
  while true do
     leftNode = prev = curr;
     $curr\rightarrownext$;
@@
        if head.id == me
           leftNode = prev = head;
@@
           end if
          continue;
        end if
       return head;
     end if
@@
           return null;
        end if
        $curr\rightarrownext$;
        if curr.id $\neq$ me
           return curr;
        end if
       leftNode = prev = curr;
        head = curr;
        $curr\rightarrownext$;
     end if
@@
        if delinkNode(prev, curr)
           Lines 1 to 1;
           continue;
       else
           leftNode == prev == head;
@@
  end if
  return null;
end
bool delinkNode(Ref prev, Ref &curr):
@@
  while currNext.mark == 1 do
@@
  end while
 temp = curr;
@@
     curr = currNext;
     return true;
  end if
 return false;
end
bool Find(Key key, Ref SH = null):
@@
     entry = registry.getByKey(key);
@@
  end if
 sId = SH.id;
  if sId == me
     leftNode = null;
     node = Search(key, SH, leftNode);
     if node == null
        return false;
     end if
    sId = node.id;
     if sId == me
        return true;
     end if
     response = Send Find(key,node) to sId;
     return response;
  end if
 response = Send Find(key,SH) to sId;
  return response;
end
@@
     entry = registry.getByKey(key);
@@
  end if
 sId = SH.id;
  if sId == me
     return InsertInSublist(key, SH);
  end if
 response = Send Insert(key,SH) to sId;
  return response;
end
bool InsertInSublist(Key key, Ref SH):
  while true do
     leftNode = null;
     rightNode = Search(key, SH, leftNode);
     if rightNode $\neq$ null
        if rightNode.id == me
           return false;
        end if
       response = Send Insert(key,rightNode) to rightNode.id;
        return response;
     end if
@@
        if SH.id == me
           continue;
        end if
@@
        return response;
     end if
@@
           continue;
        end if
     end if